L^{A}TEX Test File

Computer Science Laboratory

SRI International Menlo Park CA 94025 USA

   Rushby@csl.sri.com

Fax: +1 (415) 859-2844

If you are able to read this, you can display compressed L^{A}TEX dvi files ok.